OpenStudy (tiffany_rhodes):

A sequence X is defined recursively as follows: s1 = 100 sn+1 = 1/2 (sn + 1/3). Find lim Sn (as n approaches infinity). The formula I got for Sn = (50/ 3) * (n-1) * (1/2) ^(n-1). Also, I got the limit of Sn (as n approaches infinity) is 0. However, I'm not sure if this is correct or not.

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle s_1=100 }\) \(\color{black}{\displaystyle s_{n+1}=\frac{1}{2}\left(s_n+\frac{1}{3}\right)}\) like this?

OpenStudy (tiffany_rhodes):

yes

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle s_{n+1}=\left(\frac{3s_n+ 1}{6}\right).}\) \(\color{black}{\displaystyle s_{2}=\left(\frac{301}{6}\right).}\) \(\color{black}{\displaystyle s_{3}=\left(\frac{904}{36}\right).}\) \(\color{black}{\displaystyle s_{3}=\left(\frac{904}{6^3}\right).}\) etc, so you know intuitively.

OpenStudy (solomonzelman):

oh, excuse me, that should be 2713, not 904.

OpenStudy (solomonzelman):

We can do something like a Comparison Test, but as regards to convergence of a sequence (not series). \(\color{black}{\displaystyle s_1=1000 }\) \(\color{black}{\displaystyle s_{n+1}=\frac{2s_n}{3}}\) \(\color{black}{\displaystyle s_{n+2}=\frac{\frac{2s_n}{3}}{3}=\frac{2s_n}{3^2}}\) and then, \(\color{black}{\displaystyle s_{n+k}=\frac{\frac{2s_n}{3}}{3}=\frac{2s_n}{3^k}}\)

OpenStudy (solomonzelman):

basically, you keep dividing by 3, in the above case. and thus, it will go to 0 with more and more terms.

OpenStudy (solomonzelman):

So, for your case, where a smaller sequence is considered .... (you know)

OpenStudy (solomonzelman):

We considered a larger sequence than yours, but we can also take a smaller one. \(\color{black}{\displaystyle s_1=10 }\) \(\color{black}{\displaystyle s_{n+1}=\frac{1s_n}{2}}\) \(\color{black}{\displaystyle s_{n+2}=\frac{\frac{s_n}{2}}{2}=\frac{2s_n}{2^2}}\) and then, \(\color{black}{\displaystyle s_{n+k}=\frac{\frac{s_n}{2}}{2^{k-1}}=\frac{s_n}{2^k}}\)

OpenStudy (holsteremission):

Probably more useful would be a solution in terms of the initial condition. \[\begin{align*} s_{n+1}&=\frac{1}{2}s_n+\frac{1}{6}\\[1ex] &=\frac{1}{2}\left(\frac{1}{2}s_{n-1}+\frac{1}{6}\right)+\frac{1}{6}\\[1ex] &=\frac{1}{2^2}s_{n-1}+\frac{1}{6}\left(1+\frac{1}{2}\right)\\[1ex] &=\frac{1}{2^2}\left(\frac{1}{2}s_{n-2}+\frac{1}{6}\right)+\frac{1}{6}\left(1+\frac{1}{2}\right)\\[1ex] &=\frac{1}{2^3}s_{n-2}+\frac{1}{6}\left(1+\frac{1}{2}+\frac{1}{2^2}\right)\\[1ex] &~~\vdots\\[1ex] &=\frac{1}{2^{k+1}}s_{n-k}+\frac{1}{6}\sum_{i=0}^k\frac{1}{2^i}\\[1ex] &=\frac{1}{2^{k+1}}s_{n-k}+\frac{2-2^{-k}}{6} \end{align*}\]Set \(k=n-1\) and you get \[\begin{align*} s_{n+1}=\frac{1}{2^n}s_1+\frac{2-2^{-(n-1)}}{6}\implies s_n&=\frac{100}{2^{n-1}}+\frac{2-2^{-(n-2)}}{6}\\[1ex] s_n&=\frac{1+598\times2^{-n}}{3} \end{align*}\]Then \(\lim\limits_{n\to\infty}s_n\) is easy to find. However, it's possible (but probably unlikely), that you mean to find the limit of \(S_n\), where \(S_n\) might denote the \(n\)th partial sum of the sequence. I say unlikely because the partial sums obviously diverge.

OpenStudy (reemii):

Intuitively, ` start at initial value 100 and keep taking the average with 1/3 `.  Doing this, the distance between every new average and 1/3 is divided by 2 at each step. At the limit, the distance is 0.